A principal of N2500 is invested at 2% interest. Find the amount after 20 years if the interest is compounded (a) annually, (b) semi-annually, (c) quarterly, (d) monthly, and (e) daily. 


1
Expert's answer
2021-12-21T12:49:09-0500

"A=P(1+\frac{r}{n})^{nt}"

Compounded annually n=1


"=N2500(1+\frac{0.02}{1})^{1×20}"


"= 2500(1.02)^{20} = N3714.87"




"A=P(1+\frac{r}{n})^{nt}"

Compounded semi-annually n=2


"=N2500(1+\frac{0.02}{2})^{2×20}"


"=2500(1.01)^{40} =N3722.16"




"A=P(1+\frac{r}{n})^{nt}"

Compounded quarterly n=4


"=N2500(1+\frac{0.02}{4})^{4×20}"


"= 2500(1.005)^{80} =N3725.85"



"A=P(1+\frac{r}{n})^{nt}"

compounded monthly n =12


"=N2500(1+\frac{0.02}{12})^{12×20}"


"=2500(1.001667)^{240} =N3728.62"




"A=P(1+\frac{r}{n})^{nt}"

compounded daily n=365


"=N2500(1+\frac{0.02}{365})^{365×20}"


"=2500(1.000054795)^{7300}= N3729.53"